# Evolution of a Single Impurity Across the Superfluid–Mott insulator Transition in the Bose–Hubbard Model

### v1

@@ VI Crossing the superfluid–Mott insulator transition at fixed impurity–bath coupling
Inside the Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 18.0$), $\Delta N(R)$ becomes almost flat, with $\Delta N_{\mathrm{max}}\lesssim 0.02$ and $R(\Delta N_{\mathrm{max}})=1$ or 2, indicating that the impurity induces only a very weak, short-range density enhancement confined to its vicinity. The tiny amplitude confirms the absence of real particle recruitment: no additional bosons are bound to the impurity, and thus no bound cluster or multi-particle defect forms. Instead, the impurity remains a nearly free defect propagating in an essentially rigid background. Thus for $U_{\mathrm{ib}}/t=-8.0$ the impurity never crosses into a localized particle-defect regime, even deep inside the Mott insulator phase.
@@
These observations provide an important contrast to the repulsive case at similar impurity-bath coupling $U_{\mathrm{ib}}/t=8.0$ in MI bath, where a vacancy defect emerges once the bath becomes incompressible. For attractive $U_{\mathrm{ib}}/t=-8.0$, the impurity is energetically unable to attract and bind a full additional boson because of the large interaction gap of the Mott insulator; the local distortion remains bounded by $|\Delta N_{\mathrm{max}}|<0.02$, and no quantized particle capture or cluster formation occurs. The resulting state is therefore a weakly dressed, nearly free defect, consistent with the vanishing bath compressibility and strong on-site repulsion.
@@
Figure 14: Radial-averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ at fixed attractive coupling $U_{\mathrm{ib}}/t=-8.0$ across SF–MI transition. Main panel: Radial impurity-centered correlator $C_{\mathrm{ib}}(R)$ for different bath interactions. In the superfluid regime ($U_{\mathrm{b}}/t=8.0$–12.0), $C_{\mathrm{ib}}(R)$ shows a smooth, extended positive halo, characteristic of a mobile light polaron. As $U_{\mathrm{b}}/t$ increases, both the amplitude and range of $C_{\mathrm{ib}}(R)$ are continuously suppressed, demonstrating the compressibility-driven collapse of bath dressing. In the deep Mott-insulating regime ($U_{\mathrm{b}}/t\geq 22.0$), the correlator becomes strictly short-ranged and nearly vanishes, indicating the formation of a nearly-free defect with no particle-defect formation. Inset: On-site contrast $C_{\mathrm{ib}}(0)$ as a function of $U_{\mathrm{b}}/t$. Its monotonic decay to nearly zero confirms that, at moderate attraction ($U_{\mathrm{ib}}/t=-8.0$), the impurity never recruits one additional bath boson and does not form a localized particle-defect, in contrast to the strong-attraction case discussed elsewhere.
@@
Figure 14 further characterizes the radical-centered impurity–bath correlator $C_{\mathrm{ib}}(R)$ along the vertical cut at fixed $U_{\mathrm{ib}}/t=-8.0$. The radial-averaged correlator $C_{\mathrm{ib}}(R)$ in the main panel shows a pronounced on-site accumulation in the weakly interacting superfluid ($U_{\mathrm{b}}/t=8.0$), followed by a broad positive tail, consistent with the extended density halo seen in $\Delta N(R)$ and indicative of a light polaron with substantial bath dressing. As $U_{\mathrm{b}}/t$ increases toward the SF–MI boundary, both the amplitude and the range of $C_{\mathrm{ib}}(R)$ are continuously suppressed, reflecting the compressibility-controlled collapse of the dressing cloud: an increasingly rigid bath can no longer support long-ranged density redistribution.
@@
Once the bath enters the Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 17.0$), $C_{\mathrm{ib}}(R)$ becomes extremely short-ranged and decays to zero within one lattice spacing. The inset highlights the on-site contrast $C_{\mathrm{ib}}(0)$, which decreases monotonically with $U_{\mathrm{b}}/t$ and reaches $|C_{\mathrm{ib}}(0)|\lesssim 10^{-3}$ for $U_{\mathrm{b}}/t>22.0$. This tiny local signal confirms that, for $U_{\mathrm{ib}}/t=-8.0$, the impurity does not bind any additional bath bosons in the Mott insulator phase: no quantized particle defect is formed, and the impurity remains a weakly dressed, nearly free defect embedded in an incompressible background. This behavior stands in sharp contrast to the stronger-attraction regime ($|U_{\mathrm{ib}}|/t\gtrsim 12.0$) discussed in Sec. V.2, where a quantized particle-defect with $\Delta N_{\max}\gtrsim 1$ is stabilized in the MI phase.
@@
Figure 15: Compressibility-controlled localization of an attractive impurity across the SF–MI transition with fixed impurity-bath coupling $U_{\mathrm{ib}}/t=-8.0$. (a) Impurity winding-number estimator $\langle W_{\mathrm{imp}}^{2}\rangle$ as a function of bath interaction $U_{\mathrm{b}}/t$. In the superfluid regime ($U_{\mathrm{b}}/t<16.7$), $\langle W_{\mathrm{imp}}^{2}\rangle$ is finite, indicating a mobile light polaron. Across the SF–MI transition the winding increases even deep in the Mott insulator phase, showing that the impurity crosses over to a nearly free defect rather than a fully localized particle-defect. (b) Bath superfluid density $\rho_{\mathrm{s}}$ extracted from global winding statistics, demonstrating the loss of phase stiffness across the superfluid–Mott insulator transition. The collapse of $\rho_{\mathrm{s}}$ correlates with the suppression of compressibility-enabled dressing, while the finite $\langle W_{\mathrm{imp}}^{2}\rangle$ in panel (a) highlights the persistence of coherent impurity motion in the Mott-insulating regime.
@@
The winding estimator of the impurity and bath-superfluid density in Fig. 15 provide a complementary view of the same compressibility-controlled crossover. In the superfluid regime the impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle$ is clearly finite, consistent with a mobile light polaron dressed by an extended density halo. As $U_{\mathrm{b}}/t$ is increased and the bath superfluid density $\rho_{\mathrm{b}}$ collapses at the SF–MI transition, the winding is increases even for the largest $U_{\mathrm{b}}/t$ considered. This confirms that, for $U_{\mathrm{ib}}/t=-8.0$, the impurity evolves into a nearly free defect whose weak residual dressing is captured by the small but finite cumulative bath-density deformation in $\Delta N(R)$ and the short-ranged correlator $C_{\mathrm{ib}}(R)$.
@@ VII Summary of main results
Taken together, our results provide a unified microscopic picture of how a single impurity localizes in the two-dimensional Bose–Hubbard model when the impurity–bath coupling is attractive. Along the interaction-driven trajectory in a fixed superfluid bath ($U_{\mathrm{b}}/t=13.3$), increasing $|U_{\mathrm{ib}}|/t$ drives a continuous winding-collapse crossover from a mobile light polaron with an extended dressing cloud, to a compact heavy polaron, and finally to a tightly bound multi-particle cluster. This evolution is captured consistently by the winding number estimator of the impurity, the real-space impurity-centered correlators $C_{\mathrm{ib}}(R)$ and cumulated bath-density deformation $\Delta N(R)$, and the quasiparticle parameters $E_{\mathrm{p}}$, $m^{*}/m_{0}$, and $Z_{0}$: the winding on the impurity is suppressed, the cloud contracts, the effective mass grows, and the residue is strongly reduced while the bath remains globally superfluid.
In the Mott-insulating regime, incompressibility qualitatively reshapes the impurity physics. For sufficiently strong repulsive and attractive couplings, the impurity nucleates quantized vacancy- and particle defect, respectively: the impurity winding collapses, $\Delta N(R)$ saturates at integer values (e.g., $\Delta N_{\max}\simeq-1$ for a vacancy and $\Delta N_{\max}\gtrsim 1$ for a bound cluster), and the deformation becomes confined to within one or two lattice spacings. The impurity thus turns into a fully localized defect, and the Mott-insulating bath accommodates it by carrying an integer particle deficit or surplus in its vicinity, directly reflecting the incompressibility and interaction gap of the Mott-insulating background.
A complementary viewpoint is provided by compressibility-controlled trajectories at fixed impurity–bath coupling, exemplified here by the intermediate attraction $U_{\mathrm{ib}}/t=-8.0$. As the bath interaction $U_{\mathrm{b}}/t$ is increased from the deep-superfluid regime, through the critical region, and into the Mott insulator phase, the bath superfluid density and compressibility collapse and the extended density halo around the impurity shrinks and eventually disappears. At the same time, the impurity winding increases and the deformation amplitude $|\Delta N_{\max}|$ is smaller than 0.02, with no quantized vacancy or cluster formation. Along such trajectories, impurity localization is therefore governed by a purely compressibility-controlled suppression of coherent dressing, and the impurity crosses over from a light polaron in a compressible superfluid bath to a weakly dressed, nearly free defect in an incompressible Mott insulator. The comparison of these regimes highlights that impurity localization in the Bose–Hubbard model can proceed either via interaction-driven winding-collapse in a superfluid bath, or via a compressibility-controlled collapse of polaronic dressing when the bath is tuned across the SF–MI transition.
@@ VIII Experimental realization
Single-site–resolved imaging with quantum gas microscopes provides direct access to the real-space observables introduced in this work [18, 19, 20]. In particular, the impurity-centered correlator $C_{\mathrm{ib}}(\mathbf{R})$ can be reconstructed from conditional density maps by post-selecting snapshots in which the impurity is found at position $\mathbf{r}_{\mathrm{imp}}$ and averaging the bath occupation at $\mathbf{r}_{\mathrm{imp}}+\mathbf{R}$. Radial integration of these maps yields the cumulative deformation $\Delta N(R)$, directly visualizing the formation of depletion bubbles or bound clusters around the impurity. Moreover, repeated preparation and readout of the impurity trajectory would allow one to access winding-number statistics and thereby probe the crossover from coherent motion to localized defects.
Taken together, the combination of tunable lattice depth, controllable interspecies interactions, and single-site readout provides a realistic experimental route to explore all impurity regimes identified in our QMC study: from mobile light polarons through heavy polarons to saturated bubbles (for repulsive impurity–bath couplings) or bound clusters (for attractive couplings) in the superfluid bath, and from nearly free defects to fully localized vacancy- or particle-like defects in the Mott-insulating regime.
@@ IX Discussion and Outlook
The present work provides a unified microscopic picture of impurity localization across compressible superfluid phase and incompressible Mott-insulator phase within the two-dimensional Bose–Hubbard model. By combining worldline winding diagnostics with real-space bath deformations, we have established that the impurity localization phenomenon is in fact controlled by two distinct mechanisms depending on the bath compressibility and impurity-bath coupling.
@@
In the superfluid regime, where long-wavelength phase fluctuations are gapless and the bath remains fully compressible, the buildup of dressing is a continuous process. As impurity-bath coupling $|U_{\mathrm{ib}}|/t$ is increased at fixed $U_{\mathrm{b}}/t$, the impurity’s winding number and the associated coherent motion are gradually suppressed: the winding estimator of the impurity decreases and eventually collapses as the impurity approaches a self-trapped regime. In parallel, the cumulative bath density deformation $\Delta N(R)$ evolves smoothly from a weak, extended halo, shown by a broad, low-amplitude $C_{\mathrm{ib}}(R)$, to a compact, high-contrast structure in which the response $C_{\mathrm{ib}}(R)$ is concentrated within one or two lattice spacings. On the repulsive side, this limiting configuration corresponds to a saturated depletion bubble with $\Delta N_{\max}\simeq-1$, whereas on the attractive side it becomes a dense bound cluster with $\Delta N_{\max}\gtrsim 1$. Throughout this evolution the bath remains globally superfluid, so the localization is controlled entirely by the interaction-driven collapse of winding and dressing, rather than by any loss of phase coherence in the bath.
@@
In contrast, in the Mott-insulating regime, the situation changes qualitatively. The bath becomes incompressible, long-wavelength fluctuations are frozen, and only local particle–hole excitations with discrete energies remain. The impurity can no longer dress itself continuously; instead, localization proceeds through the formation of quantized density defects: for repulsive coupling, a single bath boson is expelled from the impurity vicinity, producing a vacancy-defect with $N_{\mathrm{min}}\!\approx\!-1$; for attractive coupling, one boson is captured to form a particle-defect with $N_{\mathrm{max}}\!\approx\!1$. These two regimes represent mirror images of each other, illustrating how the sign of impurity-bath coupling $U_{\mathrm{ib}}/t$ determines whether the impurity locally depletes or enriches the background. The jump in $\langle W_{\mathrm{imp}}^{2}\rangle$, together with the discrete jump in local bath occupancy, demonstrates that the impurity’s coherence vanishes abruptly and the underlying density response is quantized. This dual behavior embodies the interplay between quantum coherence and the integer-filling constraint of the Mott insulator state.
@@
From a broader perspective, our results suggest that a single impurity acts as a local quantum probe of the many-body gap. The correlations between the impurity winding, the extremal cumulative deformation $\Delta N_{\rm ext}$ (either $\Delta N_{\min}$ or $\Delta N_{\max}$), and the global bath compressibility provide a sensitive diagnostic of the onset of incompressibility: as the gap opens, the winding collapses and the integrated density deviation becomes quantized. This type of impurity-based diagnosis is conceptually similar in spirit to local spectroscopy and impurity thermometry in quantum gases, but is realized here entirely within an equilibrium path-integral framework. The same observables could be measured directly using site-resolved quantum gas microscopy: $C_{\mathrm{ib}}(R)$ can be extracted from conditional density maps averaged over snapshots in which the impurity occupies site $\mathbf{r}_{\mathrm{imp}}$, while the cumulative bath-density response $\Delta N(R)$ can be reconstructed from the radially averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$. Tracking the impurity trajectory over repeated experimental runs would give access to the winding-number statistics and hence to the crossover from mobile to localized regimes.
@@
Several extensions of this work naturally arise. First, finite-temperature effects may reveal the thermal activation of particle–hole excitations in the Mott-insulating background, leading to a temperature-dependent delocalization crossover. Second, by introducing multiple impurities, one can study induced interactions mediated by the bath, ranging from weak RKKY-like coupling in the superfluid to bound “defect pairs” in the Mott insulator. Third, relaxing the constraint $t_{\mathrm{a}}\!=\!t_{\mathrm{b}}$ introduces an additional tuning knob: in the limit $t_{\mathrm{a}}\!\ll\!t_{\mathrm{b}}$ the impurity becomes effectively heavy, approaching the static defect problem, whereas $t_{\mathrm{a}}\!\gg\!t_{\mathrm{b}}$ realizes a fast impurity moving through a slow, strongly correlated bath, allowing comparison with nonequilibrium polaron dynamics. Finally, by focusing on the two-impurity sector one can address bipolaron formation directly. Even in the simplest case without explicit impurity–impurity interactions, bath-mediated attraction may bind two impurities into a composite object whose binding energy and effective mass can be mapped across the $(U_{\mathrm{ib}}/t,U_{\mathrm{b}}/t)$ plane. Including a tunable direct impurity–impurity coupling would further allow one to disentangle bath-mediated and bare interactions, and remains fully accessible within the same sign-problem-free QMC framework.

### v2

@@ VI Crossing the superfluid–Mott insulator transition at fixed impurity–bath coupling
Deep in the Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 18.0$), $\Delta N(R)$ becomes almost flat, with $\Delta N_{\mathrm{max}}\lesssim 0.02$ and $R(\Delta N_{\mathrm{max}})=1$ or 2, indicating that the impurity induces only a very weak, short-range density enhancement confined to its vicinity. The tiny amplitude confirms the absence of quantized particle recruitment: no additional bosons are bound to the impurity, and thus no bound cluster or particle defect forms. Instead, the impurity remains a nearly free defect propagating in an essentially rigid background. Thus for $U_{\mathrm{ib}}/t=-8.0$ the impurity never crosses into a particle-defect regime, even deep inside the Mott-insulating phase.
@@
These observations provide an important contrast to the repulsive case at similar impurity–bath coupling $U_{\mathrm{ib}}/t=8.0$ in the MI, where a vacancy defect emerges once the bath becomes incompressible. For attractive $U_{\mathrm{ib}}/t=-8.0$, the impurity is energetically unable to attract and bind a full additional boson because of the large interaction gap of the Mott insulator; the local distortion remains bounded by $\Delta N_{\mathrm{max}}\lesssim 0.02$, and no quantized particle capture or cluster formation occurs.
@@
Figure 14: Radial-averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ at fixed attractive coupling $U_{\mathrm{ib}}/t=-8.0$ across the SF–MI transition. Main panel: Radial impurity-centered correlator $C_{\mathrm{ib}}(R)$ for different bath interactions. In the superfluid regime ($U_{\mathrm{b}}/t=8.0$–12.0), $C_{\mathrm{ib}}(R)$ shows a smooth, extended positive halo, characteristic of a mobile light polaron. As $U_{\mathrm{b}}/t$ increases, both the amplitude and range of $C_{\mathrm{ib}}(R)$ are continuously suppressed, demonstrating the compressibility-driven collapse of bath dressing. Deep in the Mott-insulating regime ($U_{\mathrm{b}}/t\geq 22.0$), the correlator becomes strictly short-ranged and nearly vanishes, indicating a nearly free defect with no particle-defect formation. Inset: On-site contrast $C_{\mathrm{ib}}(0)$ as a function of $U_{\mathrm{b}}/t$. Its monotonic decay to nearly zero confirms that, at moderate attraction ($U_{\mathrm{ib}}/t=-8.0$), the impurity never recruits one additional bath boson and does not form a self-trapped particle defect.
@@
Figure 14 further characterizes the radial-centered impurity–bath correlator $C_{\mathrm{ib}}(R)$ along the vertical cut at fixed $U_{\mathrm{ib}}/t=-8.0$. The radial-averaged correlator $C_{\mathrm{ib}}(R)$ in the main panel shows a pronounced on-site accumulation in the weakly interacting superfluid ($U_{\mathrm{b}}/t=8.0$), followed by a broad positive tail, consistent with the extended density halo seen in $\Delta N(R)$ and indicative of a light polaron with substantial bath dressing. As $U_{\mathrm{b}}/t$ increases toward the SF–MI boundary, both the amplitude and the range of $C_{\mathrm{ib}}(R)$ are continuously suppressed, reflecting the compressibility-controlled collapse of the dressing cloud: an increasingly rigid bath can no longer support long-ranged density redistribution.
@@
Once the bath enters the Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 17.0$), $C_{\mathrm{ib}}(R)$ becomes extremely short-ranged and decays to zero within one lattice spacing. The inset highlights the on-site contrast $C_{\mathrm{ib}}(0)$, which decreases monotonically with $U_{\mathrm{b}}/t$ and reaches $|C_{\mathrm{ib}}(0)|\lesssim 10^{-3}$ for $U_{\mathrm{b}}/t>22.0$. This tiny local signal confirms that, for $U_{\mathrm{ib}}/t=-8.0$, the impurity does not bind any additional bath bosons in the Mott-insulating phase: no quantized particle defect is formed, and the impurity remains a weakly dressed, nearly free defect embedded in an incompressible background. This behavior stands in sharp contrast to the strong-attraction regime ($|U_{\mathrm{ib}}|/t\gtrsim 12.0$) discussed in Sec. V.2, where a quantized particle defect with $\Delta N_{\max}\simeq 1$ is stabilized in the MI phase.
@@
Figure 15: Compressibility-controlled crossover of an attractive impurity across the SF–MI transition at fixed $U_{\mathrm{ib}}/t=-8.0$. (a) Impurity winding-number estimator $\langle W_{\mathrm{imp}}^{2}\rangle$ as a function of bath interaction $U_{\mathrm{b}}/t$. In the superfluid regime ($U_{\mathrm{b}}/t<16.7$), $\langle W_{\mathrm{imp}}^{2}\rangle$ is finite, indicating a mobile light polaron. Across the SF–MI transition the winding increases and remains finite even deep in the Mott-insulating regime, showing that the impurity crosses over to a nearly free defect rather than self-trapping into a particle defect. (b) Bath superfluid density $\rho_{\mathrm{b}}$ extracted from global winding statistics, demonstrating the loss of phase stiffness across the superfluid–Mott-insulator transition. The collapse of $\rho_{\mathrm{b}}$ coincides with the disappearance of extended dressing, while the finite $\langle W_{\mathrm{imp}}^{2}\rangle$ in panel (a) highlights the persistence of coherent impurity motion in the Mott-insulating regime.
@@
The winding estimator of the impurity and bath superfluid density in Fig. 15 provide a complementary view of the same compressibility-controlled crossover. In the superfluid regime the impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle$ is clearly finite, consistent with a mobile light polaron dressed by an extended density halo. As $U_{\mathrm{b}}/t$ is increased and the bath superfluid density $\rho_{\mathrm{b}}$ collapses at the SF–MI transition, the impurity winding increases and remains finite for the largest $U_{\mathrm{b}}/t$ considered. This confirms that, for $U_{\mathrm{ib}}/t=-8.0$, the impurity evolves into a nearly free defect whose weak residual dressing is captured by the small cumulative bath-density deformation in $\Delta N(R)$ and the short-ranged correlator $C_{\mathrm{ib}}(R)$.
@@ VII Summary of main results
Taken together, our results provide a unified microscopic picture of how a single impurity self-traps in the two-dimensional Bose–Hubbard model when the impurity–bath coupling is attractive. Along the interaction-driven trajectory in a fixed superfluid bath ($U_{\mathrm{b}}/t=13.3$), increasing $|U_{\mathrm{ib}}|/t$ drives a continuous winding-collapse crossover from a mobile light polaron with an extended dressing cloud, to a compact heavy polaron, and finally to a tightly bound multi-particle cluster.
In the Mott-insulating regime, incompressibility qualitatively reshapes the impurity physics. For sufficiently strong repulsive and attractive couplings, the impurity binds into quantized vacancy and particle defects, respectively: the impurity winding collapses, $\Delta N(R)$ saturates at integer values (e.g., $\Delta N_{\min}\simeq-1$ for a vacancy and $\Delta N_{\max}\simeq 1$ for a particle defect), and the deformation becomes confined to within one or two lattice spacings. The impurity thus turns into a fully self-trapped defect, and the Mott-insulating bath accommodates it by carrying an integer particle deficit or surplus in its vicinity, directly reflecting the incompressibility of the Mott-insulating background.
A complementary viewpoint is provided by compressibility-controlled trajectories at fixed impurity–bath coupling, exemplified here by the intermediate attraction $U_{\mathrm{ib}}/t=-8.0$. As the bath interaction $U_{\mathrm{b}}/t$ is increased from the deep-superfluid regime, through the critical region, and into the deep Mott-insulating regime, the bath superfluid stiffness collapses and the extended density halo around the impurity shrinks and eventually disappears. At the same time, the impurity winding increases and the deformation amplitude $\Delta N_{\max}$ decays below $0.02$, with no quantized particle capture. Along such trajectories, the impurity dynamics are governed by a purely compressibility-controlled suppression of extended dressing, and the impurity crosses over from a light polaron in a compressible superfluid bath to a weakly dressed, nearly free defect in an incompressible Mott insulator. The comparison of these regimes highlights that impurity self-trapping in the Bose–Hubbard model can proceed either via interaction-driven winding collapse in a superfluid bath, or via a compressibility-controlled collapse of polaronic dressing when the bath is tuned across the SF–MI transition.
@@ VIII Experimental realization
Single-site–resolved imaging with quantum gas microscopes provides direct access to the real-space observables introduced in this work [18, 19, 20]. In particular, the impurity-centered correlator $C_{\mathrm{ib}}(\mathbf{r})$ can be reconstructed from conditional density maps by post-selecting snapshots in which the impurity is found at position $\mathbf{r}_{\mathrm{imp}}$ and averaging the bath occupation at $\mathbf{r}_{\mathrm{imp}}+\mathbf{r}$. Radial integration of these maps yields the cumulative deformation $\Delta N(R)$, directly visualizing the formation of depletion bubbles or bound clusters around the impurity. Moreover, repeated preparation and readout of the impurity trajectory would allow one to access winding-number statistics and thereby probe the crossover from coherent motion to self-trapping into defect-bound states.
Taken together, the combination of tunable lattice depth, controllable interspecies interactions, and single-site readout provides a realistic experimental route to explore all impurity regimes identified in our QMC study: from mobile light polarons through heavy polarons to saturated bubbles (for repulsive impurity–bath couplings) or bound clusters (for attractive couplings) in the superfluid bath, and from nearly free defects to fully self-trapped vacancy- or particle-like defects in the Mott-insulating regime.
@@ IX Discussion and Outlook
The present work provides a unified microscopic picture of impurity self-trapping across the compressible superfluid phase and the incompressible Mott-insulating phase within the two-dimensional Bose–Hubbard model. By combining worldline winding diagnostics with real-space bath deformations, we have established that impurity self-trapping is controlled by two distinct mechanisms depending on the bath compressibility and impurity–bath coupling.
@@
In the superfluid regime, where long-wavelength phase fluctuations are gapless and the bath remains fully compressible, the buildup of dressing is a continuous process. As the impurity–bath coupling $|U_{\mathrm{ib}}|/t$ is increased at fixed $U_{\mathrm{b}}/t$, the impurity’s winding number and the associated coherent motion are gradually suppressed: the winding estimator of the impurity decreases and eventually collapses as the impurity enters a self-trapped regime. In parallel, the cumulative bath-density deformation $\Delta N(R)$ evolves smoothly from a weak, extended halo, reflected by a broad, low-amplitude $C_{\mathrm{ib}}(R)$, to a compact, high-contrast structure in which the response $C_{\mathrm{ib}}(R)$ is concentrated within one or two lattice spacings. On the repulsive side, this limiting configuration corresponds to a saturated depletion bubble with $\Delta N_{\min}\simeq-1$, whereas on the attractive side it becomes a dense bound cluster with $\Delta N_{\max}\gtrsim 1$. Throughout this evolution the bath remains globally superfluid, so the self-trapping is controlled entirely by the interaction-driven collapse of winding and dressing, rather than by any loss of phase coherence in the bath.
@@
In contrast, deep in the Mott-insulating regime, the situation changes qualitatively. The bath becomes incompressible, long-wavelength fluctuations are frozen, and only local particle–hole excitations with discrete energies remain. The impurity can no longer dress itself continuously; instead, self-trapping proceeds through defect binding with quantized density rearrangements: for repulsive coupling, a single bath boson is expelled from the impurity vicinity, producing a vacancy defect with $\Delta N_{\min}\!\approx\!-1$; for attractive coupling, one boson is captured to form a particle defect with $\Delta N_{\max}\!\approx\!1$. These two regimes represent mirror images of each other, illustrating how the sign of impurity–bath coupling $U_{\mathrm{ib}}/t$ determines whether the impurity locally depletes or enriches the background. The collapse of $\langle W_{\mathrm{imp}}^{2}\rangle$, together with the discrete change of the bath occupation, demonstrates that the impurity’s coherence vanishes abruptly and the underlying density response is quantized.
@@
From a broader perspective, our results suggest that a single impurity acts as a local quantum probe of the many-body gap. The correlations between the impurity winding, the extremal cumulative deformation $\Delta N_{\rm ext}$ (either $\Delta N_{\min}$ or $\Delta N_{\max}$), and the global bath compressibility provide a sensitive diagnostic of the onset of incompressibility: as the gap opens, extended dressing collapses and the integrated density deformation becomes quantized. This type of impurity-based diagnosis is conceptually similar in spirit to local spectroscopy and impurity thermometry in quantum gases, but is realized here entirely within an equilibrium path-integral framework. The same observables could be measured directly using site-resolved quantum gas microscopy: $C_{\mathrm{ib}}(\mathbf{r})$ can be extracted from conditional density maps averaged over snapshots in which the impurity occupies site $\mathbf{r}_{\mathrm{imp}}$, while the cumulative bath-density response $\Delta N(R)$ can be reconstructed from the radially averaged impurity-centered correlator $C_{\mathrm{ib}}(\mathbf{r})$. Tracking the impurity trajectory over repeated experimental runs would give access to winding-number statistics and hence to the crossover from mobile to self-trapped regimes.
@@
Several extensions of this work naturally arise. First, finite-temperature effects may reveal the thermal activation of particle–hole excitations in the Mott-insulating background, leading to a temperature-dependent de-self-trapping crossover. With multiple impurities, one can probe bath-mediated impurity–impurity interactions: in the superfluid they emerge as weak, long-range effective forces transmitted by compressible fluctuations, whereas in the Mott insulator they can manifest as tightly bound composite states formed via defect binding. Third, relaxing the constraint $t_{\mathrm{a}}\!=\!t_{\mathrm{b}}$ introduces an additional tuning knob: in the limit $t_{\mathrm{a}}\!\ll\!t_{\mathrm{b}}$ the impurity becomes effectively heavy, approaching the static defect problem, whereas $t_{\mathrm{a}}\!\gg\!t_{\mathrm{b}}$ realizes a fast impurity moving through a slow, strongly correlated bath. Finally, by focusing on the two-impurity sector one can address bipolaron formation directly. Even in the simplest case without explicit impurity–impurity interactions, bath-mediated attraction may bind two impurities into a composite object whose binding energy and effective mass can be mapped across the $(U_{\mathrm{ib}}/t,U_{\mathrm{b}}/t)$ plane. Including a tunable direct impurity–impurity coupling would further allow one to disentangle bath-mediated interactions and bare impurity-impurity interactions, and remains fully accessible within the same sign-problem-free QMC framework.